# Gravity versus astrophysics in black hole images and photon rings: equatorial emissions and spherically symmetric spacetimes

### v1

@@ Appendix C Additional parametric explorations
The aim of this section is to justify the choice of the relevant parameters considered in the main text and to survey a range of additional ones that go beyond the scope of this paper. All the simulations of this Appendix, unless differently stated, take into account the synchrotron pitch angle in the computation of the emissivity (21) and are performed at an observing frequency of 345 GHz. Also, we generally only consider the peaks position of the first photon ring of a disc with a mixed velocity profile around a black hole of mass $M=6.5\times 10^{9}$ $\mathrm{M}_{\odot}$ surrounded by a vertical magnetic field.
@@ C.1 Observing frequency
The results of Figures 7 and 8, computed for $\nu_{\mathrm{em}}=345$ GHz, as well as the corresponding interpretations, are very similar to the ones obtained at 230 GHz, so that it is of no interest to plot or repeat them in the case of the higher observing frequency and we will just briefly explain why this resemblance is unsurprising. The observing frequency only intervenes in the computation of the observed intensity in equation (4.2), the emitted intensity and frequency being unchanged. Thus, value of the redshift factor of equation (4.2) is modified, but not its radial dependence, so that the main impact of the observing frequency is on the value of radiative flux density received in a given band and not on the peak position of normalised intensity cuts. Beyond GR tests, we recall the importance of observing at several frequencies to get physical insights on the complex and multi-wavelength emission process and to improve the interferometric coverage for image reconstruction (Johnson et al. 2023).
C.2 Higher order photon rings
Figure 13: Positions of the $n=2$ peaks for all the possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=1.0$ (pink), a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.5$ (violet) and Schwarzschild black holes of mass $M=6.4\times 10^{9}$ $\mathrm{M}_{\odot}$ (hatched turquoise) and $M=6.3\times 10^{9}$ $\mathrm{M}_{\odot}$ (hatched teal). Critical curves are represented as dashed lines.
We simulate the second photon ring adaptively, in the $n=2$ lensing band, with a radial resolution $\mathrm{res}^{b,\>n=2}=0.001$ µas. This higher lensed observable appear to have more clearly separated intensity peaks positions given a sufficient resolving power. Nonetheless, as for the $n=1$ image, these peaks cannot be used as a discriminating probe of any conservative spacetime when the uncertainty on the mass-to-distance estimate is too high; here a precision of 3% is needed. This limited relaxation of the $M/D$ precision constraint can be understood by recalling that photon rings track the critical curve for increasing $n$ and that the position of the critical curve is a $M/D$ highly dependent quantity. The relative measurements introduced in Section 8.2 are again very degenerate even for this more strongly lensed observable, the asymmetry being of the order of 0.01% for all considered spacetimes.
@@ C.3 Magnetic field configuration
Figure 14: Positions of the $n=1$ peaks for all the possible emission configurations in the case of a Schwarzschild BH (green), a RZ BH with $a_{1}|_{\,\epsilon=0}=0.1$ (pink) and a RZ BH with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). Critical curves are represented as dashed lines and the magnetic field is radial.
@@
Figure 15: $n=1$ geodesics emitted at $r=2,3,6,15\,M$ and directions of emission (blue) as well as those of a vertical (green) and a radial (red) magnetic field. The inset shows the value of the the pitch angle $\theta_{\mathrm{B}}$ of the first lensed photons for the two, vertical and radial, magnetic field configurations in the case of a disc with mixed velocity profile.
@@
The peak positions obtained for a radial magnetic field, see Figure C.2, are equivalent in terms of astrophysical-geometrical degeneracy with respect to the ones assuming a vertical magnetic field in Figure 9. However, contrary to the vertical case, all bands in Figure C.2 are located strictly above the critical curve. Let us analyse the reasons of this different behaviour. First of all, it is worth noticing that the synchrotron pitch angle $\theta_{\mathrm{B}}$ only appears in the emissivity of equation (21) which is then the only quantity affected by a change of the magnetic field configuration. Also, from its definition, that is, the angle between the magnetic field vector and the emission direction in the rest frame of the emitter, it is clear that its value would be different for a same point source if one considers its direct or secondary image, because the directions of emission of a $n=0$ or $n=1$ photon differ. Let us then treat $n=0$ and $n=1$ separately and consider a disc seen face-on around a Schwarzschild black hole for simplicity [^32] We expect our statements to be just slightly modified when adding a small inclination or small metric deviations.. Let $\boldsymbol{k}=\boldsymbol{p}^{\mathrm{em}}/\hbar$, with $\hbar$ the reduced Planck constant, be the 4-vector tangent to the photon’s geodesic at emission. Its space-like projection $\boldsymbol{K}$ orthogonal to the 4-velocity $\boldsymbol{u}$ of the emitter, that is, in the rest-frame of the emitter, reads (Vincent et al. 2024): $\boldsymbol{K}=\boldsymbol{k}+(\boldsymbol{k}\cdot\boldsymbol{u})\,\boldsymbol{u}$. In the rest frame of the emitter, the magnetic field 4-vector is also orthogonal to the 4-velocity of the emitter: $\boldsymbol{B}\cdot\boldsymbol{u}=0$, so that $\boldsymbol{K}\cdot\boldsymbol{B}=\boldsymbol{k}\cdot\boldsymbol{B}$. Let us now define $\boldsymbol{\bar{K}}$ and $\boldsymbol{\bar{B}}$ as the unit vectors along $\boldsymbol{K}$ and $\boldsymbol{B}$ respectively. As $\boldsymbol{k}\cdot\boldsymbol{u}=-\omega$, with $\omega$ the positive pulsation of the photon as measured by the emitter, $\boldsymbol{\bar{K}}=\boldsymbol{K}/\omega$. As for $\boldsymbol{\bar{B}}$, considering that $\boldsymbol{u}=u^{t}(\boldsymbol{\partial}_{t}+\chi\boldsymbol{\partial}_{r}+\Omega\boldsymbol{\partial}_{\varphi})$, with $\chi=u^{r}/u^{t}$ and $\Omega=u^{\varphi}/u^{t}$, via equation (B7), it is straightforward to get [^33] We note that for a pure Keplerian velocity, $\chi=0$ and so $\boldsymbol{\bar{B}}_{\text{Radial}}=\mathbf{e}_{r}$.:
@@
$&\boldsymbol{\bar{B}}_{\text{Vertical}}=-\mathbf{e}_{\theta}\\ &\boldsymbol{\bar{B}}_{\text{Radial}}=\frac{1}{\sqrt{-(g_{tt}+\chi^{2}g_{rr})}}(\,\chi\sqrt{g_{rr}}\,\mathbf{e}_{t}+\sqrt{-g_{tt}}\,\mathbf{e}_{r})\>.\end{aligned}$ (C1)
It follows then
@@
$&(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})_{\text{Vertical}}=-\frac{k^{\theta}}{\omega}\sqrt{g_{\theta\theta}}\\ &(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})_{\text{Radial}}=\left(\,\frac{k^{r}}{\omega}-\chi\,\frac{k^{t}}{\omega}\right)\frac{\sqrt{-g_{tt}g_{rr}}}{\sqrt{-(g_{tt}+\chi^{2}g_{rr})}}\>.\end{aligned}$ (C2)
The synchrotron pitch angle is defined from these quantities as:

### v2

@@ Appendix C Additional parametric explorations
The aim of this section is to justify the choice of the relevant parameters considered in the main text and to survey a range of additional ones which go beyond the scope of this paper. All the simulations of this Appendix, unless differently stated, take into account the synchrotron pitch angle in the computation of the emissivity (21) and are performed at an observing frequency of 345 GHz. Also, we generally only consider the peaks position of the first photon ring of a disc with a mixed velocity profile around a black hole of mass $6.5\times 10^{9} M\mathrm{M}_{\odot}$ surrounded by a vertical magnetic field.
@@ C.1 Observing frequency
The results of Figures 7 and 8, computed for $\nu_{\mathrm{em}}=345$ GHz, as well as the corresponding interpretations are very similar to the ones obtained at 230 GHz, so it is of no interest to plot or repeat them in the case of the higher observing frequency, and we therefore only briefly explain why this resemblance is unsurprising. The observing frequency only intervenes in the computation of the observed intensity in equation (4.2), the emitted intensity and frequency being unchanged. Thus, the value of the redshift factor of equation (4.2) is modified, but not its radial dependence, so that the main impact of the observing frequency is on the value of radiative flux density received in a given band and not on the peak position of normalised intensity cuts. Beyond gravity tests, we note the importance of observing at several frequencies to get physical insights on the complex and multi-wavelength emission process and to improve the interferometric coverage for image reconstruction (Johnson et al. 2023).
C.2 Higher-order photon rings
Figure 13: Positions of the $n=2$ peaks for all the possible emission configurations in the case of black holes of mass $6.5\times 10^{9} M\mathrm{M}_{\odot}$ described by a Schwarzschild metric (green), a RZ metric with $a_{1}|_{\,\epsilon=0}=0.1$ (pink), a RZ metric with $\epsilon\,|_{\,a_{1}=0}=0.5$ (violet) and Schwarzschild black holes of mass $6.4\times 10^{9} M\mathrm{M}_{\odot}$ (hatched turquoise) and $6.3\times 10^{9} M\mathrm{M}_{\odot}$ (hatched teal). Dashed lines represent critical curves.
We simulated the second photon ring adaptively in the $n=2$ lensing band, with radial resolution $\mathrm{res}^{b,\>n=2}=0.001$ µas. This higher lensed observable appears to have more clearly separated intensity peaks positions given a sufficient resolving power. Nonetheless, as for the $n=1$ image, these peaks cannot be used as a discriminating probe of any conservative space-time when the uncertainty on the mass-to-distance estimate is too high; here a precision of 3% is needed. This limited relaxation of the $M/D$ precision constraint can be understood by recalling that photon rings track the critical curve for increasing $n$ and that the position of the critical curve is a $M/D$ highly dependent quantity. The relative measurements introduced in Section 8.2 are again very degenerate even for this more strongly lensed observable, the asymmetry being of the order of 0.01% for all considered space-times.
@@ C.3 Magnetic field configuration
Figure 14: Positions of the $n=1$ peaks for all the possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=0.1$ (pink), and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). Critical curves are represented as dashed lines and the magnetic field is radial.
@@
Figure 15: $n=1$ geodesics emitted at r=2,3,6,15Mr=2,3,6,15\,M and directions of emission (blue) as well as those of a vertical (green) and a radial (red) magnetic field. The inset shows the value of the pitch angle, $\theta_{\mathrm{B}}$, of the first lensed photons for the two, vertical and radial, magnetic field configurations in the case of a disc with mixed velocity profile.
@@
The peak positions obtained for a radial magnetic field (see Figure C.2) are equivalent in terms of astrophysical-geometrical degeneracy with respect to the ones assuming a vertical magnetic field in Figure 9. However, contrary to the vertical case, all bands in Figure C.2 are located strictly above the critical curve. We analysed the reasons of this different behaviour. First of all, it is worth noticing that the synchrotron pitch angle, $\theta_{\mathrm{B}}$, only appears in the emissivity of equation (21) which is then the only quantity affected by a change of the magnetic field configuration. Also, from its definition, that is, the angle between the magnetic field vector and the emission direction in the rest frame of the emitter, it is clear that its value would be different for a same point source if one considers its direct or secondary image, because the directions of emission of a $n=0$ or $n=1$ photon differ. We then treated $n=0$ and $n=1$ separately and considered a disc seen face-on around a Schwarzschild black hole for simplicity. [^32] We expect our statements to be just slightly modified when adding a small inclination or small metric deviations. Let $\boldsymbol{k}=\boldsymbol{p}^{\mathrm{em}}/\hbar$, with $\hbar$ the reduced Planck constant, be the 4-vector tangent to the photon’s geodesic at emission. Its space-like projection, $\boldsymbol{K}$, orthogonal to the 4-velocity of the emitter, $\boldsymbol{u}$, that is, in the rest-frame of the emitter, reads (Vincent et al. 2024) $\boldsymbol{K}=\boldsymbol{k}+(\boldsymbol{k}\cdot\boldsymbol{u})\,\boldsymbol{u}$. In the rest frame of the emitter, the magnetic field 4-vector is also orthogonal to the 4-velocity of the emitter: $\boldsymbol{B}\cdot\boldsymbol{u}=0$, so that $\boldsymbol{K}\cdot\boldsymbol{B}=\boldsymbol{k}\cdot\boldsymbol{B}$. We then defined $\boldsymbol{\bar{K}}$ and $\boldsymbol{\bar{B}}$ as the unit vectors along $\boldsymbol{K}$ and $\boldsymbol{B}$ respectively. As $\boldsymbol{k}\cdot\boldsymbol{u}=-\omega$, with $\omega$ the positive pulsation of the photon as measured by the emitter, $\boldsymbol{\bar{K}}=\boldsymbol{K}/\omega$. As for $\boldsymbol{\bar{B}}$, considering that $\boldsymbol{u}=u^{t}(\boldsymbol{\partial}_{t}+\chi\boldsymbol{\partial}_{r}+\Omega\boldsymbol{\partial}_{\varphi})$, with $\chi=u^{r}/u^{t}$ and $\Omega=u^{\varphi}/u^{t}$, via equation (B7), it is straightforward to get [^33] We note that for a pure Keplerian velocity, $\chi=0$ and so $\boldsymbol{\bar{B}}_{\text{Radial}}=\mathbf{e}_{r}$.
@@
$\begin{aligned} &\boldsymbol{\bar{B}}_{\text{Vertical}}=-\mathbf{e}_{\theta}\\ &\boldsymbol{\bar{B}}_{\text{Radial}}=\frac{1}{\sqrt{-(g_{tt}+\chi^{2}g_{rr})}}(\,\chi\sqrt{g_{rr}}\,\mathbf{e}_{t}+\sqrt{-g_{tt}}\,\mathbf{e}_{r})\>.\end{aligned}$ (C1)
It then follows that
@@
$\begin{aligned} &(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})_{\text{Vertical}}=-\frac{k^{\theta}}{\omega}\sqrt{g_{\theta\theta}}\\ &(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})_{\text{Radial}}=\left(\,\frac{k^{r}}{\omega}-\chi\,\frac{k^{t}}{\omega}\right)\frac{\sqrt{-g_{tt}g_{rr}}}{\sqrt{-(g_{tt}+\chi^{2}g_{rr})}}\>.\end{aligned}$ (C2)
The synchrotron pitch angle is defined from these quantities as